[Hallc_running] [New Logentry] Hall C RC update: Wednesday, August 10, 2022

gaskelld at jlab.org gaskelld at jlab.org
Wed Aug 10 08:40:02 EDT 2022

Logentry Text:
Beam went off yesterday at 8 am for the pass change - by 2 pm beam was back in the hall and the day shift crew had checked the beam position with the carbon hole target.  At that time, the SHMS dipole tripped several times and required intervention from Ellen and Steve L. Production data-taking for Q2=2.45, W=3.2, middle epsilon, SHMS central began around 5:30 pm.  That setting was completed, and about 3/4 of the next setting (SHMS left) when the large valve downstream of the scattering chamber closed and would not open at about 1 am.   The vacuum looks ok on both sides of the valve - EES came in to troubleshoot and concluded the issue might be with the valve itself.  This will be investigated today.  The SHMS small angle rotation is on hold until this resolved.

We will have the usual RC meeting today at 11 am . Zoom link:


RC daily update page:



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4026960
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220810/4d97bf20/attachment.html>

More information about the Hallc_running mailing list